answersLogoWhite

0

Assuming that there are an equal number of even and odd faces on the eight-sided die, then the probability of rolling an even number is simply 4 in 8, or 1 in 2, or 0.5.

User Avatar

Wiki User

∙ 15y ago

What else can I help you with?

Related Questions

What is the probability of rolling divisors of 6?

when rolling a number cube when rolling a number cube what is the probability of rolling an even number or a divisor of 10? rolling an even number or a divisor of 10?


Rolling an even number on a die?

Probability of rolling an even number on a die is 1/2.


What is the probability of rolling an even number or prime number on a number cube?

The probability of rolling an even number on a die is 3 in 6 or 1 in 2. The probability of rolling a prime on a die is 3 in 6 or 1 in 2, but one of those primes is also even. Simply add the probabilities and you find that the probability of rolling an even number or a prime on a die is 5 in 6.


What is the question of the problem if rolling a number cube labeled from 1 to 6 and the probability is one half?

There could be many questions: What is the probability of rolling an even number. What is the probability of rolling an odd number. What is the probability of rolling a number less than 4. What is the probability of rolling a number more than 3. What is the probability of rolling 1,4, or 6. Basically it could be any question about the probability of rolling half of the faces.


What is the probability of rolling a 2 on a die and then re-rolling the die and getting an even number?

The probability of rolling a 2 is 1 in 6. The probability of rolling an even number is 3 in 6. The probability of doing both, on two rolls, is 3 in 36, or 1 in 12.


What are chances of rolling an even number on a 6 sided dice?

The probability of rolling an even number is 1/2.


When rolling three number cubes what is the probability of rolling a 5a number less than 3and an even number?

The probability is 57/216 = 19/108


What is the probability of rolling an even number on a number cube?

1/2


Find the probability of rolling a 2 on a number cube and then an even number on two consecutive rolls of a number cube?

1/6,3/6 ----------------------------------------------------------------------------------------------------- The probability of rolling a 2 is: P(2) = 1/6 The probability of rolling an even number is: P(even) = 1/2 The result on the second roll is independent of the result in the first roll. The probability of rolling a 2 and then rolling an even number is: P(2,even no.) = (1/6) ∙ (1/2) = 1/12 = 0.08333... ≈ 8.33%


When rolling a number cube twice what is the probability of rolling at least one even number?

one fourth


What is the probability of rolling an even number with 3 dice?

With one roll of three dice, the probability is 7/8.


The probability of rolling an even number when number cube is rolled twice?

Because 3/6 of the sides on a number cube have even numbers, the probability of rolling even on one number cube is 1/2(equivalent of 3/6). But since you're rolling twice, you multiply the probability of one by itself (therefore rolling 2 number cubes). So: 1/2x1/2=1/4 The probability of rolling an even number when a number cube is rolled twice is 1/4, 25%, or 1 out of 4.

Trending Questions
How much should a 6 foot tall 13 year old boy weigh? Does Minnesota State University of Mankato really have a high STD rate among students in comparison with other colleges? How does the impression made by a line graph change when you make the horizontal scale shorter but keep the virtical scale the same? If you were to flip a coin eight times what is the probability of getting heads at least six times? When can the standard error of the estimate be used to construct a prediction interval about a value y'? What is the red percent on logo quiz? Can a probability be greater than 1? What is the average number of tornadoes per year in tornado alley? How many pounds are in 4000000 tons? How does logistic model of population growth differ from exponential model? Which best describes how theretical probability is determined? What is the probability of choosing a month starting with the letter m? How many different three digit numbers can be written using digits from the set 4 5 6 7 or 8 without any repeating digits? What is the probability of throwing a 5 on a dice? What is the probability of choosing a red marble from a jar containing 5 red 6 green and 4 blue marbles? Is zero is an odd number or even? What are some ways in which data collected may be inaccurate? What is the probability of drawing the 8 of hearts? How is data represented in a graph or figure? The sum of two numbers is 20 One number is three times as large as the second number What are the two numbers?